# Taking the Next Step with Generative Artificial Intelligence: The Transformative Role of Multimodal Large Language Models in Science Education

### v2

@@ Abstract.
The integration of Artificial Intelligence (AI), particularly Large Language Model (LLM)-based systems, in education has shown promise in enhancing teaching and learning experiences. However, the advent of Multimodal Large Language Models (MLLMs) like GPT-4 with vision (GPT-4V), capable of processing multimodal data including text, sound, and visual inputs, opens a new era of enriched, personalized, and interactive learning landscapes in education. Grounded in theory of multimedia learning, this paper explores the transformative role of MLLMs in central aspects of science education by presenting exemplarily powerful learning scenarios. Possible applications for MLLMs could range from content creation to tailored support for content learning, fostering competencies in scientific practices, and providing assessment and feedback. These scenarios are not limited to text-based and uni-modal formats but can be multimodal, increasing personalization, accessibility, and potential learning effectiveness. Besides its potential, challenges such as data protection and ethical considerations become more salient, calling for robust frameworks to ensure responsible integration. This paper underscores the necessity for a balanced approach in implementing MLLMs, where the technology complements rather than supplants the teacher’s role, ensuring an effective and ethical use of AI in science education. It calls for further research to explore the nuanced implications of MLLMs on the evolving role of teachers and to extend the discourse beyond science education to other disciplines. Through the exploration of potentials, challenges, and future implications, we aim to contribute to a preliminary understanding of the transformative trajectory of MLLMs in science education and beyond.
@@ 1. Introduction
Science education encompasses a broad spectrum of activities, from acquiring scientific knowledge and engaging in scientific practices to effective communication about scientific findings and ideas (Council et al.,, 2012). These aspects are designed to prepare students for the complex, multifaceted challenges they will encounter in today’s world (OECD,, 2018) and, therefore, are often part of 21st-century skills frameworks (e.g., science education: (McComas,, 2014)). The essence of science learning is inherently multimodal. It necessitates that students engage in a range of activities: reading and writing scientific texts, interpreting diagrams, drawing models, analyzing data, designing solutions, and synthesizing information across various modalities. These activities shape the understanding of scientific knowledge and the development of appropriate domain-specific skills (Lemke,, 1998; Van Leeuwen and Kress,, 1990; Doran et al.,, 2021). Additionally, it requires students to seamlessly shift between these different modes. Representations such as images, graphics, tables, technical symbols, formulas, equations, graphs, and video as well as audio data, are used in the learning process. This multimodality nature of science learning demands science educators to facilitate learning with multimodal materials and provide students opportunities to engage in multimodal activities.
”Multimodal learning” (as defined and used herein) has a cognitive foundation based on the multimedia learning theory (CTML) posited by Mayer (e.g., Mayer, 1997, 2021). Combining multimodal representations like text and images can enhance knowledge acquisition, improving the integration into a coherent mental model. Mayer, (1997); Schnotz and Bannert, (2003) have provided strong evidence for multimodal approaches in science education and state that, in accordance with Mayer, (1997) that learning with multimodal material is highly effective. The real-time generation of multimodal material offers many opportunities for adaptive learning. Therefore, multimodality is not only an integral part of science education itself but, from the perspective of cognitive science, can be harnessed to build effective learning environments.
With the latest advances in Artificial Intelligence (AI), Large Language Models (LLMs) have become increasingly popular in the general educational context cf .(Seßler et al.,, 2023; Moore et al.,, 2022; Abdelghani et al., 2023a, ) and science education cf. (Bewersdorff et al.,, 2023; Küchemann et al.,, 2023), with strong potential to advance multimodal learning. The use of LLMs in education has grown mainly due to their ability to understand and generate text data (Lee et al., 2023a, ). Since LLMs merely focus on the textual mode, they generally do not capture the variety of multimodal representations in science. With the emergence of MLLMs, we are now able to meet the necessity for multimodality in science education. Researchers are able to use MLLM, such as GPT-4V, to read scoring rubrics and grade students’ drawn models (Lee and Zhai,, 2023). MLLMs can process and create content not only in text but across multiple forms of data, including images, audio, and video (Bommasani et al.,, 2021). This shift from text-centric approaches opens a new era in science education, where the capabilities of LLMs are expanded to meet the multimodal demands of science education.
In this article, we articulate the potential of MLLMs, specifically in the field of science education. Building upon prior research on LLMs in education (e.g., (Kasneci et al.,, 2023; Zhai,, 2023; Lee et al., 2023b, ) we assess the potential of MLLMs in generally reshaping educational processes. Guided by Mayer’s CTML, we explore the potential of MLLMs in exemplary scenarios of multimodality in instructional strategies and design, enhanced student engagement, assessment, and feedback. We will also discuss the challenges related to implementing and applying MLLMs in the classroom, such as data protection and ethical concerns, which become even more pressing with multimodal systems. These multimodal systems are potentially capable of processing not only (mostly) anonymous textual data but also, depending on the specific use cases, personal data like voice recordings or handwriting images. Tackling these issues is crucial to responsibly integrating MLLMs and fully realizing their potential to improve education (Lee et al., 2023b, ).
2. Framework
@@ 2.1. Core Elements of Science Education
Science education is crucial in preparing informed citizens capable of navigating an increasingly complex world (OECD,, 2018). This preparation involves imparting a comprehensive understanding of core scientific concepts and crosscutting ideas essential for building a solid foundation in science (Council et al.,, 2012). Alongside this robust content knowledge, the development of scientific thinking and practical skills is critical. Engaging students in scientific inquiry and hands-on investigations not only cultivates their critical thinking but also equips them with practical abilities essential for scientific practices (Bybee,, 1997; Flick,, 1993). Central to science education is also the emphasis on effective communication skills. Students should be able to articulate complex scientific ideas with clarity and precision.
@@
Central tasks for teachers are the creation of engaging content and empowering students’ learning. Empowering students’ learning in science education involves not only fostering the acquisition of content knowledge but also encouraging engagement in scientific practices and the communication of scientific findings and ideas. Assessment and providing effective feedback is a third crucial task of teachers. These central aspects of science education will guide the presentation of the exemplary potentials of MLLMs in chapter 3.
2.2. Large Language Models
@@
Large Language Models (LLMs) are a subset of Artificial Intelligence (AI) focusing on processing and generating human-like text, enabling a wide range of applications from content creation to problem-solving (Brown et al.,, 2020). Those deep learning architectures are characterized by their ability to understand context, generate coherent responses, and provide textual content in real time.
@@
Incorporating human feedback through the training of reward models and applying reinforcement learning in the fine-tuning step (Ouyang et al.,, 2022) has significantly enhanced human-like language generation and led to the success of today’s popular language models by OpenAI, ChatGPT and GPT-4 (OpenAI,, 2023). The PaLM models (Chowdhery et al.,, 2022; Anil et al.,, 2023) from Google are setting a focus on a smaller and denser architecture, achieving similar performance. Also, in the open-source landscape, there are strong competitors like Meta’s LLaMA model (Touvron et al., 2023a, ; Touvron et al., 2023b, ), and the Falcon-40B architecture (Penedo et al.,, 2023). These models have sparked further development, with researchers fine-tuning them and creating open-source variants like Vicuna (Zheng et al.,, 2023) and Alpaca (Taori et al.,, 2023).
@@
The impact of these LLMs has been profound and led to novel advances across various industries. They have enabled innovative approaches in areas such as healthcare (Chintagunta et al.,, 2021; Enarvi et al.,, 2020), finance (Dowling and Lucey,, 2023), journalism (Pavlik,, 2023), and creative writing (Yuan et al.,, 2022), demonstrating the strength and versatility of LLMs. The advances in the generation of human-like texts have opened up new opportunities and challenges have in different domains (Kaddour et al.,, 2023). This includes a meaningful and constructive application in educational contexts, considering remaining issues like correct mathematical reasoning (Imani et al., 2023a, ), reliable student error detection (Bewersdorff et al.,, 2023), or possible hallucinations in the output (Ji et al.,, 2023). Research continues to address these challenges while also improving their explainability to increase trust and reliability in the model outputs (Wu et al., 2023c, ). While LLMs have significantly impacted text-based learning (e.g. (Bewersdorff et al.,, 2023; Seßler et al.,, 2023; Küchemann et al.,, 2023)), the advent of MLLMs promises to extend these benefits to include visual, auditory, and other sensory data modalities.
2.3. Multimodal Large Language Models
@@
Building upon the foundation of LLMs, MLLMs are a further advancement considering their ability to process and generate content across various data types, including text, images, audio, and video (Bommasani et al.,, 2021). These models are built to understand, interpret, and respond to multiple modalities, enriching context information beyond merely text-based capabilities. Much initial research focuses on the combination of visual and textual understanding as a multimodal input, yet they predominantly generate output in text form only (Alayrac et al.,, 2022; Zhu et al.,, 2023; Huang et al.,, 2023; Ye et al.,, 2023). To accomplish this joint understanding of textual and visual input, developers often integrate existing image encoders with pre-trained LLMs. They enable this combination by incorporating additional modules that project the different data forms into a shared embedding space (Li et al.,, 2023; Liu et al.,, 2023; Zhang et al., 2023b, ).
@@
However, the capabilities of MLLMs are not limited to images and text; they can also process combinations of text and video data (Li et al.,, 2023; Maaz et al.,, 2023). For example, this can be achieved by dissecting video into audio and a series of images, then encoding and embedding each element separately to capture the context of the video format (Zhang et al., 2023b, ). Taking a more comprehensive approach, multimodal inputs can also move beyond mere dual-modality and include additional sensor data (Driess et al.,, 2023) or speech (Chen et al.,, 2023). An example of this advancement is PandasGPT (Su et al.,, 2023), which integrates six different modalities as input for LLMs.
@@
Enhancements in multimodal capabilities, such as more advanced structures for audio processing (Zhang et al., 2023a, ), have the potential to significantly boost performance. The concept of a singular model capable of handling text, images, audio, and video for both input and output is still relatively nascent. Recently, Wu et al. published NExT-GPT as a general-purpose any-to-any open-source MLLM (Wu et al., 2023b, ) and OpenAI released GPT-4 Vision (OpenAI,, 2023), marking an important further development in this research. Also, Google announced an new LLM-based AI system they named Gemini, which is, according to Google ”built from the ground up for multimodality - reasoning seamlessly across text, images, video, audio, and code” (DeepMind,, 2023).
2.4. Adaptive, Multimodal Learning
@@
Human cognitive processing of information occurs through various channels, each handling different forms such as spoken or written text and images. The notion that combining multiple modalities can enhance knowledge acquisition is documented in Mayer’s theory of multimedia learning (1997, 2021) as well as in Schnotz & Bannert’s theory of text-picture integration (Schnotz and Bannert,, 2003). Based on cognitive models of information processing, such as Paivio’s Dual Coding Theory (1991) (Paivio,, 1991), the so-called multimedia effect is the well-supported hypothesis that active processing of both textual and pictorial information enhances integration into a coherent mental model (for a recent meta-analysis of the multimedia effect, see (Hu et al.,, 2021)). Knowledge acquisition becomes more complete and, thus, more effective when learners construct mental models from different modalities. Different representations can complement each other or even lead to a better overall understanding. For example, an image of a human skeleton or the depiction of an enzymatic reaction could be more effective and easier to understand than a written description, particularly considering the time required to process the information. However, the verbal representation of knowledge can complement the image, especially when the terms presented in text are unknown. Statements and deeper thoughts can also be represented visually, although not unambiguously.
@@
According to the theory of multimodal learning (CTML), multimodal representation enables the selection, organization, and integration of learning content into a coherent mental model (Mayer,, 2021). Learners first select relevant information from the presented material, both verbal and visual, to focus their attention on. The selected information is then organized in the learner’s mind into coherent structures, and finally, learners integrate the organized information with their existing knowledge to make sense of it. For instance, according to the signaling principle cf. (van Gog,, 2014), learners often have difficulty selecting important information in learning materials, so they need to be supported in their search to avoid unnecessary cognitive load, especially in text-image relations (for a meta-analysis of different studies on the signaling principle, see (Richter et al.,, 2016)). Previous research has shown that signaling, i.e., separate selection aids, can hinder learning, especially for learners with at least sufficient prior knowledge (e.g., (Richter and Scheiter,, 2019)).
@@
Second, learners have to organize their learning materials, i.e., text and images, according to their needs to keep their cognitive demands as low as necessary for efficient learning processes to take place (Mayer and Moreno,, 2003). Third, according to (Mayer,, 2021), the integration of multimodal representations requires generative activities on the part of the learner. It is therefore essential that multimodal representations are actively transformed into coherent mental models. As will be discussed below in relation to specific science education contexts, MLLMs offer a potential added value to this three-step process by enabling learners and teachers to adapt multimodal representations themselves.
As the Cognitive Theory of Multimedia Learning (CTML) refers to the processing of multimodal information in the context of teaching-learning processes cf. (Mayer,, 2021) it provides a suitable conceptual framework for discussing the potentials of MLLMs in (science) education.

### v3

@@ Abstract.
The integration of Artificial Intelligence (AI), particularly Large Language Model (LLM)-based systems, in education has shown promise in enhancing teaching and learning experiences. However, the advent of Multimodal Large Language Models (MLLMs) like GPT-4 with vision (GPT-4V), capable of processing multimodal data including text, sound, and visual inputs, opens a new era of enriched, personalized, and interactive learning landscapes in education. This paper derives a theoretical framework for integrating MLLMs into Multimodal Learning. This framework serves to explore the transformative role of MLLMs in central aspects of science education by presenting exemplary innovative learning scenarios. Possible applications for MLLMs range from content creation to tailored support for learning, fostering engagement in scientific practices, and providing assessments and feedback. These applications are not limited to text-based and uni-modal formats but can be multimodal, thus increasing personalization, accessibility, and potential learning effectiveness. Despite the many opportunities, challenges such as data protection and ethical considerations become salient, calling for robust frameworks to ensure responsible integration. This paper underscores the necessity for a balanced approach in implementing MLLMs, where the technology complements rather than supplants the educators’ roles, ensuring an effective and ethical use of AI in science education. It calls for further research to explore the nuanced implications of MLLMs for educators and to extend the discourse beyond science education to other disciplines. Through developing a theoretical framework for the integration of MLLMs into Multimodal Learning and exploring the associated potentials, challenges, and future implications, this paper contributes to a preliminary examination of the transformative role of MLLMs in science education and beyond.
@@ 1. Introduction
Science education encompasses a broad spectrum of activities, from acquiring scientific knowledge and engaging in scientific practices to effective communication about scientific findings and ideas (US: NRC, (2012)). These activities are essential to prepare students for the complex, multifaceted challenges they will encounter in their future lives (OECD,, 2018) and, therefore, contribute to the 21st-century competencies (e.g., Kay and Dardis, (2016); for science education: McComas, (2014)). The essence of science learning is inherently multimodal, which requires students to engage in science and engineering practices with different modalities: reading and writing scientific explanations or arguments, interpreting diagrams, drawing models, flowcharts, diagrams, analyzing and visualizing data, and designing solutions. These activities facilitate the understanding of scientific knowledge and improve domain-specific competencies (Lemke,, 1998; Van Leeuwen and Kress,, 1990; Doran et al.,, 2021). Additionally, engaging in these activities requires students to seamlessly shift between different modes. Representations such as images, graphics, tables, technical symbols, formulas, equations, process workflows, graphs, videos, and audio data are essential learning modes in the learning process. This multimodal nature of science learning demands science educators to facilitate learning with multimodal materials and provide students with opportunities to engage in multimodal activities.
Aligned with the multimodal nature of science itself, research has documented evidence regarding the effectiveness of multimodal learning for fostering learning outcomes in science education. ”Multimodal learning” (as defined and used herein) has a cognitive foundation based on the multimedia learning theory (CTML) posited by Mayer (e.g., Mayer, (1997, 2021)). Combining multimodal representations like text and images can enhance knowledge acquisition, improving hence the knowledge pieces into a coherent multi-faceted mental model. In alignment with findings from (Mayer,, 1997), Schnotz and Bannert, (2003) provided strong evidence that multimodal approaches are effective in promoting (science) learning. This is because, human beings learn via different information channels and each channel can bear certain amount of working load. Learners struggle when information is intensified and provided in a single modality. Multimodal learning offers opportunities for learners to reduce the cognitive load for each channel. Despite this potential, researchers found that students seldom have the opportunity to engage in high-quality multimodal learning due to the shortage of materials that teachers could generate and provide.
With the latest advances in Artificial Intelligence (AI), specifically in generative AI, Multimodal Large Language Models (MLLMs) have shown great potential to address this gap. MLLMs are a subset of AI models focusing on processing and generating human-like text, images, videos, audio, etc., enabling thus a wide range of applications from content creation to problem-solving (Bommasani et al.,, 2021). They can also transform unstructured data into structured content and vice versa (Borisov et al., 2022b, ). MLLMs such as GPT-4 or Gemini enable educators to generate learning resources (e.g., lesson plans (Lee et al., 2024)), understand scoring rubrics, and assign scores to students’ written responses or drawn models, meeting the needs of multimodality in science education (Lee and Zhai,, 2023).
In this article, we articulate the potentials and challenges of MLLMs in reshaping the science learning processes and environments. Building upon Mayer’s CTML, we first develop an AI-enhanced multimodal learning framework and then present a series of exemplary scenarios where MLLMs are used to increase the modalities of learning, including instructional strategies and design, student engagement, assessment, and feedback. We also demonstrate the capacity of MLLMs to provide adaptive multimodal learning support to students with diverse needs. At last, we discuss the challenges related to implementing and applying MLLMs in the classroom, such as data protection and ethical concerns, which have become increasingly pressing with MLLMs
2. Review of the literature
@@ 2.1. Core Elements of Science Education
Science education is crucial in preparing informed citizens capable of navigating an increasingly complex world (OECD,, 2018). This education involves imparting a comprehensive understanding of core scientific concepts and crosscutting ideas essential for building a solid foundation in science (NRC,, 2012). Alongside this robust content knowledge, the development of scientific thinking and practical skills is critical. Engaging students in scientific inquiry and hands-on investigations not only cultivates their critical thinking but also equips them with practical abilities essential for scientific practices (Bybee,, 1997; Flick,, 1993). Moreover, such an approach may also foster their creativity, enabling them to devise effective solutions for ill-structured problems (Hathcock et al.,, 2015). Central to science education is also the emphasis on effective communication skills. Students should be able to articulate and communicate complex scientific ideas with clarity and precision (NRC,, 2012).
@@
Central tasks for educators are creating engaging content and empowering students’ learning. In science education, the latter involves not only fostering the acquisition of content knowledge but also encouraging engagement in scientific practices and the communication of scientific findings, insights, and ideas. Assessment and providing effective feedback are crucial tasks for educators. These three central aspects of science education will guide the presentation of the exemplary potentials of MLLMs in Chapter 3.
2.2. Large Generative AI Models
@@
Large Language Models. LLMs are a subset of AI models focusing on processing and generating human-like text, enabling thus a wide range of applications from content creation to problem-solving (Brown et al.,, 2020). Those deep learning architectures are characterized by their capacity to process and interpret context, generate coherent responses, and provide textual content in real time.
@@
Incorporating human feedback through the training of reward models and applying reinforcement learning in the fine-tuning step (Ouyang et al.,, 2022; Lee et al., 2023b, ) has significantly enhanced the generation of human-like language that is aligned with social norms and goals and has led to the success of today’s popular language models by OpenAI, ChatGPT and GPT-4 (OpenAI,, 2023). The PaLM models (Chowdhery et al.,, 2022; Anil et al.,, 2023) from Google are setting a focus on a smaller and denser architecture, achieving similar performance. Also, in the open-source landscape, there are strong competitors like Meta’s LLaMA model (Touvron et al., 2023a, ; Touvron et al., 2023b, ), and the Falcon-40B architecture (Penedo et al.,, 2023). These models have sparked further development, with researchers fine-tuning them and creating open-source variants like Vicuna (Zheng et al.,, 2023) and Alpaca (Taori et al.,, 2023).
@@
The impact of these LLMs has been profound and led to novel advances across various industries. They have enabled innovative approaches in areas such as healthcare (Chintagunta et al.,, 2021; Enarvi et al.,, 2020), finance (Dowling and Lucey,, 2023), journalism (Pavlik,, 2023), creative writing (Yuan et al.,, 2022), and even in scientific discovery (Romera-Paredes et al.,, 2023), demonstrating the strength and versatility of LLMs. The advances in the generation of human-like texts have opened up new opportunities and challenges in different domains (Kaddour et al.,, 2023). This includes a meaningful and constructive application in educational contexts, considering remaining issues like correct mathematical reasoning (Imani et al., 2023a, ), or possible hallucinations in the output (Ji et al., 2023a, ). Research continues to address these challenges while also improving their explainability to increase trust and reliability in the model outputs (Wu et al., 2023c, ). While LLMs have the potential to significantly impact text-based learning (e.g., Bewersdorff et al., 2023a ; Seßler et al., (2023); Küchemann et al., (2023)), the advent of MLLMs promises to extend these benefits to include visual, auditory, and other sensory data modalities.
@@
Multimodal Large Language Models. Building upon the foundation of LLMs, MLLMs are a further advancement considering their ability to process and generate content across various data types, including text, images, audio, and video (Bommasani et al.,, 2021). However, these models are built to understand, interpret, and respond to multiple modalities, enriching context information beyond merely text-based capabilities. Much initial research focuses on the combination of visual and textual understanding as a multimodal input, yet they predominantly generate output in text form only (Alayrac et al.,, 2022; Zhu et al.,, 2023; Huang et al.,, 2023; Ye et al.,, 2023). To accomplish this joint understanding of textual and visual input, developers often integrate existing image encoders with pre-trained LLMs. They enable this combination by incorporating additional modules that project the different data forms into a shared embedding space (Li et al., 2023b, ; Liu et al.,, 2023; Zhang et al., 2023c, ).
@@
However, the capabilities of MLLMs are not limited to images and text; they can also process combinations of text and video data (Li et al., 2023b, ; Maaz et al.,, 2023). For example, this can be achieved by dissecting video into audio and a series of images, then encoding and embedding each element separately to capture the context of the video format (Zhang et al., 2023c, ). Taking a more comprehensive approach, multimodal inputs can also move beyond mere dual-modality and include additional sensor data (Driess et al.,, 2023) or speech (Chen et al.,, 2023). An example of this advancement is PandasGPT (Su et al.,, 2023), which integrates six different modalities as input for LLMs.
@@
Enhancements in multimodal capabilities, such as more advanced structures for audio processing (Zhang et al., 2023a, ), have the potential to significantly boost performance. The concept of a singular model capable of handling text, images, audio, and video for both input and output is still relatively nascent. Recently, Wu et al. published NExT-GPT as a general-purpose any-to-any open-source MLLM (Wu et al., 2023b, ) and OpenAI released GPT-4 Vision (OpenAI,, 2023) and GPT-4 Turbo [^1] https://chatopenai.de/gpt-4-turbo/, marking an important further development in this research. Also, Google announced an new LLM-based AI system they named Gemini, which is, according to Google ”built from the ground up for multimodality - reasoning seamlessly across text, images, video, audio, and code” (DeepMind,, 2023). This kind of reasoning in combination with other types of deep learning and explainability approaches (Wu et al.,, 2020; Borisov et al., 2022a, ; Rombach et al.,, 2022) has the potential to advance science (Wong et al.,, 2023) and education. Often, these AI systems are labeled Multimodal Foundation Models (Li et al., 2023a, ). Following Fu et al., (2023), we will use the term ’Multimodal Large Language Models’ in our paper.
2.3. Adaptive Multimodal Learning
@@
Information processing occurs through various channels, each handling different forms of information, such as spoken or written language or images. The dual nature of information (verbal vs. non-verbal) and the associated (specific) processing of different information types is called multimodality. The empirically well-supported idea that combining multiple modalities can enhance knowledge acquisition is documented in Mayer’s theory of multimedia learning (Mayer,, 1997, 2021) as well as in Schnotz & Author’s theory of text-picture integration (Schnotz and Bannert,, 2003). Based on cognitive models of information processing, such as Paivio’s Dual Coding Theory (Paivio,, 1991), the so-called multimedia effect is the well-supported hypothesis that active processing of both verbal (textual) and non-verbal (visual) information enhances integration into a coherent mental model (for a recent meta-analysis of the multimedia effect, see Hu et al., (2021)). Knowledge acquisition becomes more complete and less ambiguous and, thus, more effective when learners construct mental models from different modalities. Different representations can complement each other or even lead to a better overall understanding. For example, an image of a human skeleton or the depiction of an enzymatic reaction could be more effective and easier to understand than a written description, particularly when considering the time required to process the information. In summary, combining textual and pictorial information can effectively support learning. This can be further specified by the Cognitive Theory of Multimedia Learning (CTML).
@@
According to CTML, multimodal representation enables the selection, organization, and integration of learning content into a coherent mental model (Mayer,, 2021). Learners first select relevant information from the presented material to focus their attention on. The selected information is then organized in the learner’s mind into coherent structures, and finally, learners integrate the organized information with their existing knowledge to make sense of it. For instance, according to the signaling principle (cf. van Gog, (2014)), learners often have difficulty selecting important information in learning materials, so they need to be supported in their search to avoid unnecessary cognitive load, especially in text-image relations (for a meta-analysis of different studies on the signaling principle, see Richter et al., (2016)). Previous research has shown that signaling, i.e., separate selection aids, can hinder learning, especially for learners with at least sufficient prior knowledge (e.g., Richter and Scheiter, (2019)).
@@
Second, learners have to organize their learning materials, i.e., text and images, according to their needs to keep their cognitive demands as low as necessary for efficient learning processes to take place (Mayer and Moreno,, 2003). Third, according to Mayer, (2021), the integration of multimodal representations requires generative activities on the part of the learner. It is, therefore, essential that multimodal representations are actively transformed into coherent mental models by the student.
If we summarise the processes of selection, organization, and integration and discuss them in the context of MLLM, the potential for adaptivity becomes clear. In multimedia presentations, visual and textual information is usually statically embedded in a media presentation (e.g., a textbook or a graph). However, neither text nor visual information can be easily adapted to the learner’s needs in real-time. This changes with MLLMs as they allow the learner to convert textual information into images and vice versa with simple prompts. For example, it is possible to convert text from MLLMs into mind maps, graphical illustrations, or to extract terms from a complex representation of scientific processes, making it much easier for the learner to select, organise and integrate information into complete mental models. In other words, MLLM offer a promising way of ”representational scaffolding” (Fischer et al.,, 2022) to extract terms from complex scientific representations, helping learners to select and organize, i.e., the adaptation of representational modalities to learners’ needs.
As will be discussed below in relation to specific science education contexts, MLLMs offer a potential added value to this three-step process by enabling learners and educators to adapt multimodal representations themselves. As the CTML refers to the processing of multimodal information in the context of teaching-learning processes (cf. (Mayer,, 2021)), it provides a suitable conceptual framework for the following discussion of the potential of MLLMs in (science) education.
3. Framework of Integrating MLLM into Multimodal Learning
Rooted in the Theory of Multimedia Learning (cf. 2.3) and considering the capabilities of MLLMs (cf. 2.2), we propose a framework that integrates MLLMs into multimodal learning. Positioned between the verbal (textual) and non-verbal (visual) channels (refer to Section 2.3 and Dual Coding Theory (Paivio,, 1991)), the MLLM operates as an essential processing unit that bridges these channels. In our framework, the MLLM processes visual or textual input based on the user’s requirements for adaptiveness and personalization, producing textual and/or visual outputs. The MLLM has two key functionalities (cf. Figure 1):
  1. A.
    Transforming content from text to image or vice versa: MLLMs facilitate the shift between textual and visual representations. For instance, in science education, converting tabular data into visual diagrams can provide further insights. In Figure 1, this example would be represented by the path from textual input through the MLLM to the visual output.
  2. B.
    Shifting from uni- to multimodality by adding a modality: As each channel has a limited capacity for processing information at any given time, adding a modality —supplementing text with visuals or vice versa—aims to reduce this cognitive load. Knowledge acquisition becomes more complete, less ambiguous, and thus more effective when learners construct mental models using various modalities (see 2.3). In science education, one example could be to augment the structure of an insect’s eye given in an illustration by textual information. In Figure 1, this example would be represented by both the path from the visual input through the MLLM to the textual output as well as the path from the visual input right to the visual output.
Figure 1. Framework of Integrating MLLM into Multimodal Learning
Depending on the setting, the user of the MLLM can be either the educator or the learner itself. With the MLLM, the user can generate adaptive and personalized multimodal representations. When the educator operates the MLLM, they must provide information such as competencies, student needs, and individual difficulties to the MLLM. If the student operates the MLLM on their own, information to ensure adaptivity and personalization must be provided beforehand by the educator, the student themselves, or the MLLM may indirectly retrieve this information through interaction. Adaptability ensures that the transformation and augmentation of modalities are tailored to meet the learner’s needs and competencies. It is crucial for the MLLMs to inherently apply effective multimedia design principles (Mayer,, 2024), e.g., arising from the signaling principle (van Gog,, 2014). This proposed framework represents an initial draft, laying the groundwork for further development and refinement in integrating multimodal learning with advanced capabilities of MLLMs.